OpenStudy (gamenerd123):

What is the common ratio for the geometric sequence? −15,−9,−27/5,−81/25,.

OpenStudy (tkhunny):

Have you considered dividing successive terms?

OpenStudy (gamenerd123):

I'm not sure how to do that

OpenStudy (tkhunny):

You are not sure how to divide? I doubt that. \(\dfrac{-9}{-15} = \dfrac{9}{15} = \dfrac{3}{5}\) Okay, try the next pair.

OpenStudy (gamenerd123):

Oh, so I divide the numbers?

OpenStudy (gamenerd123):

It seems that the geometric term for the two fractions are 5 and 3

OpenStudy (gamenerd123):

Idk what to do after that

OpenStudy (tkhunny):

You need a common ration. Start with the first number. Apply the ratio to get the second number. Apply the ratio again to get the third number. and etc...

OpenStudy (gamenerd123):

Well I would how to do it from 15 to 9, but I'm not sure about the fractions

OpenStudy (tkhunny):

\(15\cdot\dfrac{3}{5} = 9\)

OpenStudy (gamenerd123):

Wait, how did you get to 9?

OpenStudy (tkhunny):

You WILL be required to simplify fractions on your exam. If all else fails, just do it on your calculator. \(15\cdot\dfrac{3}{5} = \dfrac{15}{5}\cdot 3 = 3\cdot 3 = 9\)

OpenStudy (gamenerd123):

So the common ratio is 9?

OpenStudy (gamenerd123):

I mean 3?

OpenStudy (tkhunny):

?? Where did the 5 go?